A pile of thirty-nine coins contains only dimes and quarters. The value of all the coins is $7.65. If each dime was exchanged for a nickel and each quarter exchanged for a half-dollar, what would the new value, in dollars, of the coins? Thanks for your help.
Answer by TimothyLamb(4379)

x = number of dimes
y = number of quarters
---
x + y = 39
10x + 25y = 765
---
put the system of linear equations into standard form
---
x + y = 39
10x + 25y = 765
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = number of dimes = 14
y = number of quarters = 25
---
answer:
x = number of dimes = 14*0.05 = $0.70
y = number of quarters = 25*0.50 = $12.50
0.70 + 12.50 = $13.20
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---